OpenStudy (anonymous):

help please. medal. the time t required to drive a certain distance varies inversely with the speed, r. if it takes 4 hours to drive the distance at 40 miles per hour, how long will it take to drive the same distance at 55 miles per hour? a. about 2.91 hours b. about 160.00 c. 22.00 hours d. about 5.50 hour

OpenStudy (welshfella):

inverse variation :- t = k/r where k is a constant

OpenStudy (welshfella):

4 hours at 40 mph so 4 = k/40 solve this for k

OpenStudy (welshfella):

replace k in the equation by the value for k you just found this will give you the relation between t and r To find t when r = 55 plug in r = 55 and calculate t

OpenStudy (welshfella):

can you solve 4 = k/40 ?